# Visual SLAM with 3D Gaussian Primitives and Depth Priors Enabling Novel View Synthesis

### v1

@@ Abstract
Conventional geometry-based SLAM systems lack dense 3D reconstruction capabilities since their data association usually relies on feature correspondences. Additionally, learning-based SLAM systems often fall short in terms of real-time performance and accuracy. Balancing real-time performance with dense 3D reconstruction capabilities is a challenging problem. In this paper, we propose a real-time RGB-D SLAM system that incorporates a novel view synthesis technique, 3D Gaussian Splatting, for 3D scene representation and pose estimation. This technique leverages the real-time rendering performance of 3D Gaussian Splatting with rasterization and allows for differentiable optimization in real time through CUDA implementation. We also enable mesh reconstruction from 3D Gaussians for explicit dense 3D reconstruction. To estimate accurate camera poses, we utilize a rotation-translation decoupled strategy with inverse optimization. This involves iteratively updating both in several iterations through gradient-based optimization. This process includes differentiably rendering RGB, depth, and silhouette maps and updating the camera parameters to minimize a combined loss of photometric loss, depth geometry loss, and visibility loss, given the existing 3D Gaussian map. However, 3D Gaussian Splatting (3DGS) struggles to accurately represent surfaces due to the multi-view inconsistency of 3D Gaussians, which can lead to reduced accuracy in both camera pose estimation and scene reconstruction. To address this, we utilize depth priors as additional regularization to enforce geometric constraints, thereby improving the accuracy of both pose estimation and 3D reconstruction. We also provide extensive experimental results on public benchmark datasets to demonstrate the effectiveness of our proposed methods in terms of pose accuracy, geometric accuracy, and rendering performance.
@@ I INTRODUCTION
Current research, highlighted by the comprehensive investigations of Ming et al.  [ming2024benchmarking] in NeRF-based SLAM systems. They found that traditional SLAM faces challenges in matching sensor observations with 3D landmark points, making geometry estimation difficult for unobserved regions and potentially leading to tracking loss. Implicit neural representations like NeRF overcome this limitation by utilizing synthesizing new views, making them promising for visual SLAM systems. In addition, research fields including Edge computing and Reinforcement learning also integrate well on SLAM. Edge computing  [li2023predictive, zong2022cocktail] involves processing data near the source of data generation, such as on the robot itself or on a nearby device in the network. Reinforcement learning  [10406020, li2023red] can be used to develop optimal navigation strategies that allow the robot to efficiently explore unknown areas while avoiding obstacles. By learning from its environment, the robot can make decisions that balance exploration and exploitation. These approaches can complement SLAM by providing quick, localized processing while still allowing for cloud-based analysis if needed.
They categorize NeRF-based methods for SLAM into RGB-D SLAM, monocular SLAM, and LiDAR SLAM. Additionally, they provide an overview of the NeRF-based SLAM pipeline, illustrating the framework and summarizing the various approaches used in NeRF-based SLAM methods which is detailed in Figure 2. Figure 2 directly summarizes the current methods of applying NeRF in SLAM. Finally, they introduce relevant datasets and compare the results when evaluating NeRF-based SLAM methods.
@@
In this paper, we aim to leverage the fast rendering performance and promising explicit 3D scene representation of 3DGS [kerbl3Dgaussians], along with the accurate pose estimation strategy from Stereo-NEC [wang2024stereo, song2024eta], to generate dense 3D reconstructions while maintaining accurate camera pose estimation. We implement our pipeline using SplaTAM [spatam] as our foundation.
@@ II-B Neural Radiance Fields
The core principle of NeRF involves modeling a 3D scene using neural networks, encoding the scene within the weights of these networks. In practice, it utilizes multi-layer perceptrons (MLPs) to map spatial locations to the implicit surface representation of target objects or scenes. Additionally, the original NeRF approach emphasizes volume density.
@@
3D Gaussian Splatting is another technique that allows for novel view synthesis and 3D reconstruction. Unlike NeRF, which uses implicit representation for the environment, 3D Gaussian Splatting employs 3D Gaussians as primitives for explicit 3D scene representation. This rendering process enables quick, differentiable rendering of 3D scenes by projecting these 3D Gaussians onto a pixel-based image plane to obtain a rendered image. Each Gaussian is represented by Equation 1 and is parameterized by only eight values: three for its RGB color $c$, three for its center position $\boldsymbol{\mu}\in\mathbb{R}^{3}$, one for its radius $r$, and one for its opacity $o\in[0,1]$. Since 3D Gaussian splatting is differentiable, it allows for the updating of the eight parameters using gradient-based optimization by minimizing the difference between the rendered image and the input image.
@@
where $f_{i}(\cdot)$ is a function defined in Equation 1, and $c_{i}$ denotes the color of the $i$-th 3D Gaussian.
@@ III Proposed Approach
Our method inspired by Stereo-NEC [wang2024stereo], SplaTAM [spatam], and BundledSLAM [10503743] aims to provide a SLAM system with dense reconstruction using depth priors. During the tracking stage, we adopt the decoupled rotation and translation estimation strategy from Stereo-NEC, using inverse optimization. Additionally, we incorporate the 3DGS concept from SplaTAM for 3D scene representation, which enables the ability to synthesize novel views. Multi-camera SLAM systems have many advantages, quite directly resulting from the ability to incorporate information coming from a larger field of view—leading to higher robustness and better localization accuracy. The work describes an important extension and improvement of the state-of-the-art stereo SLAM system, called ORB-SLAM2, aiming at even higher precision. To achieve this, first, measurements from all cameras are projected to a common virtual camera, BundledFrame. The virtual camera is properly designed to facilitate hassle-free handling of multi-camera setups and an opportunity to properly combine data obtained from individual cameras. The benchmarks include an extensive comparison with ground truth measurements made on the state of the art SLAM system. The assessment is performed on the EuRoC datasets to evaluate the performance of the system critically. The results demonstrate the superior accuracy of the BundledSLAM system compared to existing approaches. Specifically, building on Stereo-NEC, SplaTAM, and BundledSLAM, we separately estimate rotation and translation using inverse optimization. This approach directly employs differentiable rendering and gradient-based optimization to refine both the camera pose and the parameters of the 3D Gaussian map, based on the set of online camera poses estimated so far.
@@ III-B Pose Estimation
$\textit{D}(\textbf{p})=\sum_{i=1}^{\text{n}}d_{i}f_{i}(\textbf{p})\prod_{j=1}^{i-1}(1-f_{j}(\textbf{p}))$ (3)
@@
$\textit{S}(\textbf{p})=\sum_{i=1}^{\text{n}}f_{i}(\textbf{p})\prod_{j=1}^{i-1}(1-f_{j}(\textbf{p}))$ (4)
@@ III-C Scene Optimization
In this section, we aim to reconstruct a dense, explicit volumetric representation using a set of 3D Gaussians and then update the scene through 3D Gaussian Splatting. As camera poses are updated and new Gaussians are added, there may be duplicate 3D Gaussians that need to be removed. Additionally, optimizing the parameters of the 3D Gaussians is essential for accurate scene representation. Therefore, our scene optimization process includes both 3D Gaussian parameter optimization and duplication removal.
@@ IV-B Experiment Results
Our method achieved localization accuracy with an average ATE ranging from 3.49 cm to 12.52 cm across five challenging scenarios, demonstrating its capability to deliver centimeter-level precision in difficult indoor environments. Absolute trajectory error (ATE) is a standard metric for evaluating localization accuracy in benchmarking SLAM system’s tracking performance. This high level of accuracy is highlighted in the average ATE column of Table I.
We also evaluate the novel view synthesis performance of our method using PSNR, with values ranging from 19.63 to 25.11, indicating competitive performance. Additionally, we assess geometric accuracy by computing the average depth map RMSE, which showed depth accuracy ranging from 2.81 cm to 6 cm. These strong performances in PSNR and depth map accuracy are highlighted in the average PSNR and average Depth RMSE columns of Table I, respectively.
Qualitative Evaluation: In addition to the quantitative evaluation, we also provide a qualitative assessment. As shown in Fig 4, our rendered results for both RGB and depth images on the Freiburg2_xyz sequence are compared with the input images displayed in the top row. We also visualize the silhouette map to identify new observation areas in the current frame where new 3D Gaussians can be added to the map. Specifically, the white areas indicate the regions where new 3D Gaussians will be added.
@@ V CONCLUSIONS
In this paper, we introduce an innovative RGB-D SLAM system that utilizes 3D Gaussian Splatting (3DGS) for accurate point cloud generation, enabling dense reconstruction and real-time pose estimation. Specifically, our system leverages the computational efficiency of CUDA-enabled rasterization for rendering both rgb and depth images and optimizes both camera poses and 3D Gaussians by minimizing the difference between rendered and input images. Additionally, through a rotation-translation decoupled strategy, our method successfully meets the demanding requirements of real-time performance and precise scene reconstruction.

### v2

@@ Abstract
Conventional geometry-based SLAM systems lack dense 3D reconstruction capabilities since their data association usually relies on feature correspondences. Additionally, learning-based SLAM systems often fall short in terms of real-time performance and accuracy. Balancing real-time performance with dense 3D reconstruction capabilities is a challenging problem. In this paper, we present a real-time RGB-D SLAM system that uses 3D Gaussian Splatting for both 3D scene representation and pose estimation. This method enables efficient real-time rendering and differentiable optimization through CUDA, offering an edge over NeRF. We also enable mesh reconstruction from 3D Gaussians for explicit dense 3D reconstruction. To estimate accurate camera poses, we utilize a rotation-translation decoupled strategy with inverse optimization. This involves iteratively updating both in several iterations through gradient-based optimization. This process includes differentiably rendering RGB, depth, and silhouette maps and updating the camera parameters to minimize a combined loss of photometric loss, depth geometry loss, and visibility loss, given the existing 3D Gaussian map. However, 3D Gaussian Splatting (3DGS) struggles to accurately represent surfaces due to the multi-view inconsistency of 3D Gaussians, which can lead to reduced accuracy in both camera pose estimation and scene reconstruction. To address this, we utilize depth priors as additional regularization to enforce geometric constraints, thereby improving the accuracy of both pose estimation and 3D reconstruction. We also provide extensive experimental results on public benchmark datasets to demonstrate the effectiveness of our proposed methods in terms of pose accuracy, geometric accuracy, and rendering performance.
@@ I INTRODUCTION
Current research, highlighted by the comprehensive investigations of Ming et al.  [ming2024benchmarking] in NeRF-based SLAM systems. They found that traditional SLAM faces challenges in matching sensor observations with 3D landmark points, making geometry estimation difficult for unobserved regions and potentially leading to tracking loss. Implicit neural representations, like NeRF, use multi-layer perceptrons (MLPs) to represent the scene. In addition, research fields including Edge computing and Reinforcement learning also integrate well on SLAM. Edge computing  [li2023predictive, zong2022cocktail] involves processing data near the source of data generation, such as on the robot itself or on a nearby device in the network. Reinforcement learning  [10406020, li2023red] can be used to develop optimal navigation strategies that allow the robot to efficiently explore unknown areas while avoiding obstacles. By learning from its environment, the robot can make decisions that balance exploration and exploitation. These approaches can complement SLAM by providing quick, localized processing while still allowing for cloud-based analysis if needed.
They categorize NeRF-based methods in the SLAM domain into monocular SLAM, RGB-D SLAM, LiDAR SLAM, and multi-sensor SLAM. Additionally, they provide an overview of the NeRF-based SLAM pipeline, illustrating the framework and summarizing the various approaches used in NeRF-based SLAM methods which is detailed in Figure 2. Figure 2 directly summarizes the current methods of applying NeRF in SLAM. Finally, they benchmark NeRF in the SLAM domain and provide comprehensive results to evaluate NeRF-based SLAM methods.
@@
In this paper, we aim to leverage the fast rendering performance and promising explicit 3D scene representation of 3DGS [kerbl3Dgaussians], along with the accurate pose estimation strategy from Stereo-NEC [wang2024stereo], to generate dense 3D reconstructions while maintaining accurate camera pose estimation [song2024eta]. We implement our pipeline using SplaTAM [spatam] as our foundation. The key contributions of our proposed method are as follows:
  • We employ a rotation-translation decoupled strategy combined with inverse optimization to estimate accurate camera poses.
  • We utilize 3D Gaussian primitives to represent scenes, facilitating dense reconstruction.
  • We utilize 3D Gaussian Splatting with rasterization, allowing for real-time differentiable optimization via CUDA.
  • We incorporate depth priors as additional regularization to enforce geometric constraints, thereby enhancing the accuracy of both pose estimation and 3D reconstruction.
@@ II-B Neural Radiance Fields
NeRF’s core concept revolves around modeling a 3D scene through neural networks, where the scene is encoded within the network’ parameters. In practice, it utilizes multi-layer perceptrons (MLPs) to map spatial locations to an implicit representation of the surfaces of target objects or scenes. Moreover, the vanilla NeRF approach places a strong emphasis on volume density.
@@
3D Gaussian Splatting is another technique that allows for novel view synthesis and 3D reconstruction. Unlike NeRF, which uses implicit representation for the environment, 3D Gaussian Splatting employs 3D Gaussians as primitives for explicit 3D scene representation. This process allows for fast, differentiable rendering of 3D scenes by projecting the 3D Gaussians onto a pixel-based image plane, resulting in a rendered image. Each Gaussian is represented by Equation 1 and is parameterized by only eight values: three for its RGB color $c$, three for its center position $\boldsymbol{\mu}\in\mathbb{R}^{3}$, one for its radius $r$, and one for its opacity $o\in[0,1]$. Since 3D Gaussian splatting is differentiable, it allows for the updating of the eight parameters using gradient-based optimization by minimizing the difference between the rendered image and the input image.
@@
here, $f_{i}(\cdot)$ refers to the function defined in Equation 1, and $c_{i}$ represents the color of the $i$-th 3D Gaussian.
@@ III Proposed Approach
Our method inspired by Stereo-NEC [wang2024stereo], SplaTAM [spatam], and BundledSLAM [10503743] aims to provide a SLAM system with dense reconstruction using depth priors. During the tracking stage, we adopt the decoupled rotation and translation estimation strategy from Stereo-NEC, using inverse optimization. Additionally, we incorporate the 3DGS concept from SplaTAM for 3D scene representation, which enables the ability to synthesize novel views. Multi-camera SLAM systems have many advantages, quite directly resulting from the ability to incorporate information coming from a larger field of view—leading to higher robustness and better localization accuracy. This work presents a significant extension and enhancement of the state-of-the-art stereo SLAM system, ORB-SLAM2, with the goal of achieving even greater precision. To achieve this, first, measurements from all cameras are projected to a common virtual camera, BundledFrame. The virtual camera is properly designed to facilitate hassle-free handling of multi-camera setups and an opportunity to properly combine data obtained from individual cameras. The benchmarks include an extensive comparison with ground truth measurements made on the state of the art SLAM system. The assessment is performed on the EuRoC datasets to evaluate the performance of the system critically. The results demonstrate the superior accuracy of the BundledSLAM system compared to existing approaches. Specifically, building on Stereo-NEC, SplaTAM, and BundledSLAM, we separately estimate rotation and translation using inverse optimization. This approach directly employs differentiable rendering and gradient-based optimization to refine both the camera pose and the parameters of the 3D Gaussian map, based on the set of online camera poses estimated so far.
@@ III-B Pose Estimation
$\textit{D}(\textbf{p})=\sum_{i=1}^{\text{n}}d_{i}f_{i}(\textbf{p})\prod_{k=1}^{i-1}(1-f_{k}(\textbf{p}))$ (3)
@@
$\textit{S}(\textbf{p})=\sum_{i=1}^{\text{n}}f_{i}(\textbf{p})\prod_{k=1}^{i-1}(1-f_{k}(\textbf{p}))$ (4)
@@
TABLE II: Quantitative results for evaluating depth ($\downarrow$), Structural Similarity Index Measure (SSIM $\uparrow$), and Learned Perceptual Image Patch Similarity (LPIPS $\downarrow$).
Sequence Average Depth Loss Average MS-SSIM Average LPIPS
freiburg1_desk 3.61 cm 0.846 0.244
freiburg1_desk2 3.24 cm 0.839 0.253
freiburg1_room 3.54 cm 0.810 0.252
freiburg2_xyz 2.81 cm 0.955 0.073
freiburg3_long_office_household 6.00 cm 0.857 0.204
@@ III-C Scene Optimization
In this section, we aim to reconstruct a dense representation by a set of explicit 3D Gaussians and then update the scene through 3D Gaussian Splatting. As camera poses are updated and new Gaussians are added, there may be duplicate 3D Gaussians that need to be removed. Additionally, optimizing the parameters of the 3D Gaussians is essential for accurate scene representation. Therefore, our scene optimization process includes both 3D Gaussian parameter optimization and duplication removal.
@@ IV-B Experiment Results
Figure 6: Qualitative rendering results on freiburg1 desk. The top row, from left to right, shows the input RGB, input depth map, and rasterized silhouette. The bottom row, from left to right, displays the rendered RGB, rendered depth map, and visualization of the L1 loss on the depth map.
Figure 7: Qualitative rendering results on freiburg1 room. The top row, from left to right, shows the input RGB, input depth map, and rasterized silhouette. The bottom row, from left to right, displays the rendered RGB, rendered depth map, and visualization of the L1 loss on the depth map.
Our method achieved localization accuracy with an average ATE ranging from 3.49 cm to 12.52 cm across five challenging scenarios, demonstrating its capability to deliver centimeter-level precision in difficult indoor environments. Absolute trajectory error (ATE) is a widely used metric for assessing localization accuracy when benchmarking the tracking performance of SLAM systems. The superior accuracy of our approach is evident in the average ATE values presented in Table II.
We also evaluate the novel view synthesis performance of our method using PSNR, with values ranging from 19.63 to 25.11, indicating competitive performance. Additionally, we assess geometric accuracy by computing the average depth map RMSE, which showed depth accuracy ranging from 2.81 cm to 6 cm. These strong performances in PSNR and depth map accuracy are highlighted in the average PSNR and average Depth RMSE columns of Table II, respectively.
Qualitative Evaluation: Alongside the quantitative evaluation, we offer a qualitative assessment as well. As illustrated in Fig 4, we compare our rendered RGB and depth images on the Freiburg2_xyz sequence with the input images shown in the top row. Additionally, we visualize the silhouette map to highlight areas of new observations in the current frame where new 3D Gaussians will be added to the map. Specifically, the white regions indicate where these new 3D Gaussians will be incorporated.
@@ V CONCLUSIONS
In this paper, we introduce an innovative RGB-D SLAM system that utilizes 3D Gaussian Splatting (3DGS) for accurate point cloud generation, enabling dense reconstruction and real-time pose estimation. Our system takes advantage of the computational efficiency of CUDA-enabled rasterization to render both RGB and depth images. It optimizes camera poses and 3D Gaussians by minimizing the discrepancy between the rendered images and the input images. Additionally, through a rotation-translation decoupled strategy, our method successfully meets the demanding requirements of real-time performance and precise scene reconstruction.